The average weight of Bob, John and Danny is 70 lbs. If Bob and Danny weigh 145 lbs total, how much does John weigh?
Help show work and expalin plz

Answer by JoelSchwartz(130)

(b+j+d)/3=70
The sum of bob, john and Danny is equal to two hundred and ten. Two hundred and ten divided by three is equal to seventy. The reason i divided by three is a average is the sum of all the numbers divided by the number of numbers.
b+d=145
(145+j)/3=70
145/3+j/3=70
j=65
210/3=70